OpenStudy (anonymous):

How would you keep going from here ? -4i²+16ij-4j²

OpenStudy (anonymous):

i think that is simplified . can't go any further i don't think. unless its a foil/factoring problem

OpenStudy (anonymous):

It was a foil/factoring problem at first, then I got to that, then i'm stuck but that's not the answer though :|

OpenStudy (anonymous):

okay gimme a sec to work it out

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Can I have the original question?

OpenStudy (anonymous):

4(-i^2+4I+1j^2) . . might need to get someone to double check that

OpenStudy (anonymous):

Let V = -4i+j and W = i-4j Solve: V x W

OpenStudy (anonymous):

i not a math genious sowweh i dont think im right :/

OpenStudy (anonymous):

what is that number/letter in the middle? is that 41 or 4l ?

OpenStudy (anonymous):

doing it, give me a sec.

OpenStudy (anonymous):

Original equation: Let V = -4i+j and W = i-4j Solve: V x W

OpenStudy (anonymous):

Ok !

OpenStudy (anonymous):

-4x^2 + 16ji -4j^2

OpenStudy (anonymous):

17 my bad

OpenStudy (anonymous):

17 in the middle ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how ?

OpenStudy (anonymous):

i got : -4i²+16ij-4j² though O_O

OpenStudy (anonymous):

(-4+j) (i-4j) = -4i^2 + 16ji + ji -4j^2

OpenStudy (anonymous):

yea, that's what i got earlier :D

OpenStudy (anonymous):

but i dont know how to keep going from there ?

OpenStudy (anonymous):

the answer is -8 btw

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

I don't understand how you can get 8 from that.

OpenStudy (anonymous):

Are you sure you're reading the answer correctly?

OpenStudy (anonymous):

Yes, it's -8

OpenStudy (anonymous):

That's just not possible. If the question is just asking V x W, there's nothing more you can do.

OpenStudy (anonymous):

ahh i know what's wrong now lol, it's not multiplying it's dealing with something called Dot Product

OpenStudy (anonymous):

Sorry everyone :|

hero (hero):

Well, you should post the full instructions next time

OpenStudy (anonymous):

well the instruction did not say Dot Product, but it just simply V dot W, like a little multiplying dot in the middle, so i thought it's multiply =.=

OpenStudy (anonymous):

But my fault :|

hero (hero):

Like I said, you should post the question you were working on to begin with. What you have posted originally is not really a question that can be properly addressed.